OpenStudy (lexy148):

A sugar cookie recipe that makes 3 dozen cookies requires 2 1/2cups of powered sugar. if you want to make enough cookies so that each of 30 students in you child's class gets exactly two cookies, how much powdered sugar will you need? (can someone help me to put it into a proportion)

OpenStudy (anonymous):

For each student to get 2 each you will need 60 cookies 3*12 = 36 cookies (I am not using bakery dozen) 2.5 cups of sugar will make only 36 cookies I would think you need a total of 5 dozen (5 *12 = 60) you agree thus far?

OpenStudy (anonymous):

so now lets look at how much sugar per dozen of cookie 5/2 * 1/3 = 5/6 cup of sugar per dozen 5/6 * 5 = 25/6 = 4 1/6 cups of sugar neede

OpenStudy (anonymous):

5/2 is another way to write 2 1/2 also if I need to divide by 3 same as multiplying by 1/3

OpenStudy (lexy148):

thank you i get it now

OpenStudy (anonymous):

ok